### v1

# The Southern Photometrical Local Universe Survey (S-PLUS): searching for metal-poor dwarf galaxies

Abstract
The metal content of a galaxy’s interstellar medium reflects the interplay between different evolutionary processes such as feedback from massive stars and the accretion of gas from the intergalactic medium. Despite the expected abundance of low-luminosity galaxies, the low-mass and low-metallicity regime remains relatively understudied. Since the properties of their interstellar medium resemble those of early galaxies, identifying such objects in the Local Universe is crucial to understand the early stages of galaxy evolution. We used the DR3 catalog of the Southern Photometric Local Universe Survey (S-PLUS) to select low-metallicity dwarf galaxy candidates based on color selection criteria typical of metal-poor, star-forming, low-mass systems. The final sample contains approximately 50 candidates. Spectral energy distribution fitting of the 12 S-PLUS bands reveals that 90% of the candidates are best fit by models with very low stellar metallicities. We obtained long-slit observations with the Gemini Multi-Object Spectrograph to follow-up a pilot sample and confirm whether these galaxies have low metallicities. We find oxygen abundances in the range 12 + log(O/H) (5% to 17% of the solar value), confirming their metal-poor nature. Most targets are outliers in the mass-metallicity relation, i.e. they display a low metal content relative to their observed stellar masses. In some cases, perturbed optical morphologies might give evidence of dwarf-dwarf interactions or mergers. These results suggest that the low oxygen abundances may be associated with an external event causing the accretion of metal-poor gas, which dilutes the oxygen abundance in these systems.
1 Introduction
The metallicity of the interstellar medium (ISM) is a fundamental tracer of the evolutionary process of a galaxy (Lequeux et al., 1979; Tremonti et al., 2004). It depends on several factors, such as the star formation rate and the gas-mass fraction, but it also reflects the interplay between stellar feedback and the accretion of metal-poor gas from the intergalactic medium (Lilly et al., 2013; Peng & Maiolino, 2014), driven by the processes that drive and regulate the baryon cycle within galaxies (see Maiolino & Mannucci, 2019, for a review).
Low-luminosity galaxies are less chemically evolved than more massive systems, likely due to a less efficient star formation activity and a higher fraction of metals lost in supernova events (Guseva et al., 2009; Kirby et al., 2013; McQuinn et al., 2015; Gonçalves, 2019). However, despite the expected large population of low-luminosity galaxies, the low-mass and low-metallicity regime is still relatively little studied among the galactic populations. In the standard cosmological scenario, dwarf galaxies play a crucial role in the hierarchical formation of structures, since the first objects to form and contribute, through mergers, to the growth of more massive systems were low-mass and metal-poor galaxies (Bovill & Ricotti, 2009; Bromm & Yoshida, 2011). Atek et al. (2024) have recently reported the discovery of eight faint galaxies in the redshift range between 6 and 8 with the James Webb Space Telescope. These galaxies exhibit stellar masses between and M⊙ and have oxygen abundances of 12 + log(O/H) 7, approximately 2% the solar value111In this work we adopt a solar oxygen abundance of 12 + log(O/H)⊙ = 8.69 0.04 (Asplund et al., 2021).. Nearby gas-rich, low-metallicity dwarfs are expected to be analogous to such primitive galaxies both because of their low abundances of heavy elements and the presence of intense bursts of star formation (Bouwens et al., 2016). However, due to their low intrinsic luminosity, detailed investigations of these objects at great distances is extremely complex. Therefore, it is important to combine the most detailed observations of the closest analogous objects with those of younger and more distant galaxies.
Dwarf galaxies in the Local Universe can be studied in different galaxy environments, from large under-dense regions – the voids – to rich clusters (Grossi, 2019). Dwarf galaxies in low-density regions often show peculiar gas morphology and kinematics, likely associated with external gas accretion, and they can experience strong bursts of star formation (Hoyle et al., 2012). The most metal-poor star-forming dwarfs in the Local Universe are found in such large underdense regions (Sánchez Almeida et al., 2016a).
Progress in identifying new metal-poor systems has been relatively slow. Detecting these faint galaxies requires that they are relatively nearby, and/or that they are experiencing a recent star formation episode. Robust abundance determinations are challenging to derive for faint galaxies, and because these systems are harder to find due to their low masses, the current metallicity relations are substantially underpopulated at the low-luminosity/low-mass end. Recent studies targeting large areas of the sky only managed to detect a small number of metal-poor dwarf galaxies with metallicity below 1/10th solar in the Local Universe (; Izotov et al., 2012; Hsyu et al., 2018; McQuinn et al., 2020; Lin et al., 2023). A few extremely metal poor (XMP) objects have been found at low redshifts with oxygen abundances around or slightly below 12 + log(O/H) (Izotov et al., 2009; Hirschauer et al., 2016; Izotov et al., 2018, 2019; Kojima et al., 2020; Izotov et al., 2024). These galaxies are characterized by having low stellar mass M M⊙ and very compact structures. Discovering and studying XMP galaxies is fundamental to understand the process of star formation in a metal poor ISM and to provide constraints for theoretical models of chemical evolution of low-mass galaxies.
Metal-poor galaxies can be identified using selection criteria based on either spectroscopic (Izotov et al., 2012; Guseva et al., 2017; Hirschauer et al., 2022; Zou et al., 2024) or photometric diagnostics (James et al., 2015, 2017; Hsyu et al., 2018; Kojima et al., 2020; Lin et al., 2023). Wide-area photometric surveys can provide a more efficient tool in finding low-metallicity targets, since they allow to preselect a larger number of candidates that can be subsequently confirmed or rejected with a targeted spectroscopic campaign. James et al. (2017) identified low metallicity systems in the Sloan Digital Sky Survey (SDSS) using a combination of color cuts. Oxygen abundances obtained via the direct method varied within the range 7.45 12 + log (O/H) 8.0, and about 20% of the subset of galaxies with spectroscopic observations had less than 10% of the solar value. Hsyu et al. (2018, hereafter H18) applied color selection criteria typical of low-metallicity star-forming dwarf galaxies, such as IZw18 or Leo P to galaxies selected in the SDSS. Follow-up spectroscopic studies revealed that approximately 45% of the galaxies selected using these color cuts had oxygen abundances below 1/10th of the solar value, and only 6% had lower values, % Z⊙. Lin et al. (2023, hereafter L23) performed a similar study in the Dark Energy Survey (DES) using FUV+ color cuts. The oxygen abundances of the spectroscopic subsample were found to vary between 7.4 12+log(O/H) 8.2 with a mean of 12+log(O/H) = 7.8 ( 0.13 Z⊙).
ID | S-PLUS ID | RA | DEC | zspec | log(M∗) | ||
[AB mag] | [AB mag] | [M⊙] | |||||
1 | DR4_3_SPLUS-s27s07_0011404 | 00:37:13.1 | -35:08:48 | 16.790.01 | 16.79 0.03 | 0.0138 | 8.500.11 |
2 | DR4_3_SPLUS-n14s01_0028263 | 10:00:27.6 | -17:28:25 | 17.540.02 | 17.51 0.04 | 0.0162 | 8.400.11 |
3 | DR4_3_SPLUS-n13s01_0024584 | 10:01:16.3 | -16:13:49 | 17.760.03 | 17.80 0.06 | 0.0075 | 7.530.11 |
4 | DR4_3_SPLUS-n02n27_0037849a | 12:23:23.3 | 01:48:52 | 17.070.02 | 17.00 0.04 | 0.0067 | 7.880.11 |
5 | DR4_3_SPLUS-s29s37_0046039 | 04:05:20.4 | -36:48:59 | 16.290.01 | 16.26 0.02 | 0.0027b | 7.340.11 |
6 | DR4_3_SPLUS-n03s22_0021537c | 11:57:12.0 | -02:41:12 | 17.100.02 | 17.20 0.04 | 0.0046d | 7.270.11 |
The Southern Photometric Local Universe Survey222https://splus.cloud/ (S-PLUS) is a 12-band optical survey conducted using the T80-South 0.83-m robotic telescope (T80S) located at Cerro Tololo Interamerican Observatory (Mendes de Oliveira et al., 2019). S-PLUS uses the Javalambre 12-band magnitude system (Cenarro et al., 2019), which includes four SDSS-like broad-band, a Javalambre intermediate-band filter, and seven narrow-band filters centered on prominent stellar spectral features such as [O ii]3727,3729, Caii H+K, H, H, and the Caii triplet. The panoramic camera features a single charge-coupled device (CCD) with 9.2k 9.2k pixels, a field of view of , and a pixel scale of 055 pix-1 (Benitez et al., 2014). The survey will cover an area of 9300 deg2 of the southern sky. Although there are other ongoing wide-area photometric surveys in the southern hemisphere, the S-PLUS unique 12 multi-band filter system enables a better sampling of the wavelength range between 3000 and 10000 to perform detailed analysis of the spectral energy distribution and stellar population of galaxies. Specifically, the availability of the and J0660333This narrow-band filter is centered at 6614 and is suitable for investigating the H emission line. filter imaging over a large-area of the sky makes S-PLUS a unique survey for studying star forming galaxies.
Our aim is to start building a sample of low-metallicity dwarf galaxy candidates in the southern hemisphere using the S-PLUS survey and to perform a spectroscopic follow-up observations to confirm their chemically unevolved nature. Once we have constrained their abundances we can derive scaling relations and study their global properties in order to understand whether their low metallicity are due to either a slow secular evolution or to some external events related to their local environment. Here we present the preliminary results of a pilot spectroscopic follow-up campaign of the first metal-poor candidates with the Gemini Multi-Object Spectrograph (GMOS) attached to the 8-m Gemini South telescope (Gemini-S).
This article is organized as follows. In Section 2 we present the sample selection, in Section 3 we describe the Gemini-S/GMOS observations and the data reduction. We discuss our results in Section 4, and we present our Summary and Conclusion in Section 5. In this work, we adopt the cosmological parameters = 70 km s-1 Mpc-1, = 0.3, and .
2 Sample Selection
The search for low-metallicity objects in the S-PLUS survey was performed on the third data release (DR3) which covers an area of 2000 deg2. The photometric 5 depths range from 19.1 to 20.5 mag depending on the filter (see Almeida-Fernandes et al., 2022, for details).
2.1 color cuts
Following H18, we applied color selection criteria typical of low-metallicity star-forming dwarf galaxies, such as IZw18 or Leo P. These color cuts were defined from the analysis of galaxies in the SDSS survey to identify low redshift ( 0.01–0.02), blue compact sources with colors consistent with H ii regions:
where is the uncertainty of the band magnitude, and is the filter covering the H+[N ii] lines. The fourth condition is defined to include also objects with a poorly constrained -band magnitude (H18). Taking advantage of the S-PLUS multiband filter set, we add an additional constraint, , to ensure that galaxies exhibit detectable H emission444We note that, given the characteristics of the narrow-band J0660 filter, the H emission can only be detected within the redshift range (Logroño-García et al., 2019; Mendes de Oliveira et al., 2019). (Gutiérrez-Soto et al., 2020). We applied these criteria to the S-PLUS DR3 catalog including only objects with mag ( 200000 objects). This selection returned a total of 380 candidate objects. We visually inspected the S-PLUS images to reject spurious candidates such as stars or star-forming regions located in the spiral arms of larger galaxies. After applying an absolute magnitude cut of mag555To derive the absolute magnitudes at this stage of the selection, we used the photometric redshifts derived by Lima et al. (2022) for the S-PLUS survey., which defines dwarf galaxies (Thuan & Martin, 1981; Meyer et al., 2014; Venhola et al., 2019), we obtained a final sample of 47 metal-poor dwarf galaxy candidates (see Fig. 1 as an example).
2.2 Spectral energy distribution (SED) fitting
As a first test to confirm the metal-poor nature of our candidates, we performed SED fitting of the 12 S-PLUS bands with the Code Investigating GALaxy Emission (CIGALE, Boquien et al., 2019). CIGALE combines a stellar SED with dust attenuation and emission components. The code fits the spectrum (from 3000 to 10000) of each galaxy and derives global properties such as star formation rate, stellar mass, bolometric luminosity, the minimum intensity value of the stellar radiation field. The stellar component is described by providing the relative contribution of both the young and the old stellar components to the total emission of the galaxy. Once the predicted theoretical SEDs are modeled, they are fitted to the observed SED. The quality of the fit is assessed by the best (and a reduced best defined as , with being the number of data points, Boquien et al., 2019).
CIGALE allows for a number of star formation histories (SFH) such as exponentially declined or delayed. We used a delayed exponential star formation history with e-folding times varying between and 7 Gyr for the older stellar populations and an exponential burst with e-folding time Myr for the latest star formation event. We left the stellar metallicity as a free parameter, ranging between 1/200th Z⊙ and 1 Z⊙, adopting the Calzetti et al. (2000) attenuation law. The majority of the candidates (43 out of 47) have SEDs that are best-fitted by stellar metallicity of Z = 0.0004 (1/50th Solar) or Z = 0.0001 (1/200th Solar). The stellar masses obtained from population synthesis models range between 107 and few times 109 M⊙, and they are comparable to those obtained using the relation between color and mass-to-light ratio derived by stellar population synthesis models (Bell et al., 2003)
where is the galaxy color corrected for the Milky Way extinction, absolute magnitude in AB mag units, also corrected for Milky Way extinction, and a Kroupa (2001) initial mass function (IMF) is assumed (see Table 1). This calibration has been used to derive the stellar masses of other samples of low-mass galaxies (e.g. H18, ). Therefore, we will use it throughout this work to facilitate comparison with other sets of low-metallicity dwarf galaxies (see Sections 2.3, 4.5, and 4.6). Following Bell et al. (2003) we adopt a systematic uncertainty of 25% for the mass-to-light ratios obtained with this method.
Four objects were selected to be observed with Gemini-S/GMOS between 2022 and 2023 (ID 1 - 4 in Table 1). DR4_3_SPLUS-s29s37_0046039 (ID 5 in Table 1) has oxygen abundances derived with the direct method in the literature (see Table 1). Another object – DR4_3_SPLUS-n03s22_0021537 (ID 6 in Table 1) – has emission line fluxes available from the SDSS. Its metallicity was estimated by Sánchez Almeida et al. (2016b), although not with the direct method. Therefore we downloaded the SDSS spectrum and calculated the metallicity using the same approach as for the four galaxies observed with GMOS. We display the images of these six targets in Fig. 1. The images are taken from the DESI Legacy Imaging Surveys666https://www.legacysurvey.org/ (Dey et al., 2019). Being deeper than S-PLUS, Legacy images allow us to better identify the morphology and structure of our targets.
In Fig. 2 we display examples of the SED fitting for two of our candidates.
2.3 Comparison samples
Throughout this work, we compare our targets with metal-poor dwarf galaxies selected using similar criteria in other photometric surveys such as the SDSS (H18), and the DES (L23). The only difference between the selection method between L23 and H18 is that band photometry is not available over the DES area, therefore, they modified the color-cut criterion into a condition based on the GALEX NUV filter (Martin et al., 2005). Most targets from these photometric selections have been followed up spectroscopically. For comparison, we include only galaxies with oxygen abundances derived with the direct method. Stellar masses for both samples have been recalculated using Eq. 1, assuming a Kroupa IMF.
3 Gemini-S/GMOS observations and data reduction
The observations were carried out via two Fast Turnaround runs in 2022 and 2023 with the GMOS instrument (Hook et al., 2004) in the long slit mode on the Gemini-S telescope (Program IDs GS-2022B-FT-210, GS-2023A-FT-204). We used the B600 grism with a 2′′ slit width. The wavelength range varies between 4300 and 7500 Å with a central wavelength 5900 Å. This choice allowed us to include all the major emission lines needed to determine the oxygen abundance with the direct method, including the singly ionized ionic population of oxygen traced by the [O ii]7320,7330 lines. The spectral resolution estimated as the full-width-at-half-maximum (FWHM) of the emission lines of comparison lamps at 5900 is 9 , corresponding to = 195 km s-1 ( = FWHM/2.355). The total integration time ranged between 1350 and 7200 seconds. Data were binned by 2 in the spatial direction and binned by 2 in the spectral direction only for the first galaxy observed in the 2022 observing run (ID 1). The central position of the long slit was defined according to the peak of H emission in the S-PLUS continuum-subtracted J0660 images of each galaxy. The position angle of the slit for the first target (ID 1) was set at the parallactic angle. For the remaining targets, however, the slit was oriented to cover most of the galactic disc and the H ii regions detected in the J0660 narrow-band images (Fig. 3, top panels). Spectra of CuAr arc lamps were obtained for wavelength calibrations, along with bias frames and dome flats to correct for the detector bias level and pixel-to-pixel variations, respectively. The spectrophotometric standard LTT 2415 was observed with the same experimental set-up and it was used to calibrate the targets in relative flux.
For the two GMOS runs we selected the targets that were best observable in semesters 2022B, 2023A and that were characterized by the lowest best-fit metallicities in our SED fitting analysis (see §2.2). The data were reduced with the Data Reduction for Astronomy from Gemini Observatory North and South (DRAGONS) package v3.1.0 (Labrie et al., 2023) using the standard recipe for GMOS long-slit reductions, that includes bias correction, flat-fielding, sky subtraction, wavelength calibration, flux calibration, and 1D spectra extraction.
Our aim is to derive the oxygen abundances using the direct method, which utilizes the flux ratio of auroral [O iii]4363 to strong lines such as [O iii]4959,5007, to measure the electron temperature of the gas, . Combining hydrogen recombination lines and strong oxygen emission features of [O ii] and [O iii] ions, the total oxygen abundance can be derived (Izotov et al., 2006; Osterbrock & Ferland, 2006).
4 Results
1 | 2 | 3 | 4 | |||||
4340 H | 35.25 1.31 | 43.031.60 | 40.59 0.46 | 44.000.50 | 32.06 0.89 | 37.18 1.03 | 38.22 0.90 | 41.95 0.99 |
4363 [O iii] | 12.41 1.29 | 15.001.56 | 9.09 0.30 | 9.810.32 | 5.45 0.69 | 6.28 0.80 | 4.88 0.56 | 5.33 0.61 |
4471 He i | – | – | 2.99 0.26 | 3.170.28 | – | – | – | – |
4861 H | 100.00 1.15 | 100.000.32 | 100.00 0.47 | 100.000.28 | 100.00 1.28 | 100.00 0.50 | 100.00 0.92 | 100.00 0.51 |
4959 [O iii] | 151.77 2.17 | 146.892.10 | 193.80 1.10 | 191.251.09 | 115.20 1.41 | 112.44 1.38 | 115.07 1.44 | 113.33 1.42 |
5007 [O iii] | 439.86 5.42 | 419.125.16 | 587.48 2.94 | 576.112.88 | 332.54 3.34 | 320.80 3.22 | 344.84 3.57 | 337.14 3.49 |
5876 He i | 12.65 0.66 | 10.860.57 | 11.59 0.15 | 10.900.14 | 12.53 0.47 | 11.18 0.42 | 11.22 0.37 | 10.45 0.34 |
6300 [O i] | – | – | 3.10 0.15 | 2.810.14 | – | – | – | – |
6312 [S iii] | – | – | 1.87 0.13 | 1.690.12 | – | – | – | – |
6363 [O i] | – | – | 1.22 0.15 | 1.090.13 | – | – | – | – |
6563 H | 359.70 4.25 | 265.843.14 | 311.99 1.57 | 276.031.39 | 337.31 3.31 | 269.40 2.64 | 316.20 3.22 | 274.50 2.79 |
6584 [N ii] | 11.33 0.42 | 8.340.31 | 6.98 0.15 | 6.160.13 | 7.80 0.44 | 6.21 0.35 | 8.95 0.35 | 7.75 0.30 |
6678 He i | 3.10 2.89 | 2.242.09 | 3.05 0.14 | 2.670.12 | 4.38 0.42 | 3.43 0.33 | 2.96 0.26 | 2.54 0.22 |
6716 [S ii] | 31.13 3.65 | 22.312.61 | 15.53 0.22 | 13.570.19 | 24.06 0.53 | 18.78 0.41 | 17.82 0.43 | 15.25 0.37 |
6731 [S ii] | 21.60 3.37 | 15.442.40 | 11.15 0.20 | 9.730.17 | 17.29 0.48 | 13.47 0.37 | 11.50 0.36 | 9.83 0.31 |
7065 He i | – | – | 2.45 0.11 | 2.090.09 | 2.55 0.28 | 1.90 0.21 | 2.22 0.19 | 1.85 0.16 |
7136 [Ar iii] | 5.27 0.35 | 3.500.23 | 6.50 0.10 | 5.500.08 | 2.66 0.31 | 1.96 0.23 | 6.99 0.25 | 5.77 0.21 |
7320 [O ii] | 3.48 0.35 | 2.230.22 | 2.49 0.11 | 2.080.09 | 2.29 0.30 | 1.65 0.22 | 2.27 0.25 | 1.84 0.20 |
7330 [O ii] | 3.08 0.35 | 1.970.22 | 2.26 0.11 | 1.890.09 | 2.28 0.30 | 1.64 0.22 | 1.69 0.24 | 1.37 0.19 |
2.64 | 4.84 | 0.76 | 1.94 | |||||
mean |
a in units of erg s-1 cm-2.
All the galaxies present two or three star-forming knots within the slit as can be seen in the bottom panels of Fig. 3, showing an expanded view of the 2D spectra around the H emission line. The H emission of our targets is structured in velocity along the slit.
4.1 Emission line fluxes and intensities
The apertures for the extraction of the spectra to measure the oxygen abundances were defined based on the extension of the [O iii]4363 emitting region, which is usually much more compact than H. [O iii]4363 is detected only in correspondence of the brightest H ii region of each galaxy. We defined an aperture size of 2′′ for all galaxies, selected along the spatial axis, as indicated by the gray lines in the bottom panels of Fig. 3.
In Fig. 4 we display the 1D spectra obtained for the four objects, with the prominent emission lines marked. The brightest lines are H and H recombination lines, along with the [O iii]4959,5007 lines. As can be seen in the figure, even with long exposures on a 8-meter telescope, the detection of the auroral line is challenging in these galaxies. We measured the centers, the FWHM, the equivalent widths (EW), and the fluxes of the detected emission lines by performing Gaussian fitting of the line profiles with the PySpecKit (Ginsburg et al., 2022) python package.
The emission line fluxes were corrected for internal extinction using the Balmer decrement. The Balmer decrement remains roughly constant for typical gas conditions in star-forming galaxies (Osterbrock & Ferland, 2006), and we assumed the theoretical values for Case B recombination and 104 K, [(H)/(H)]th = 2.86 and [(H)/(H)]th = 0.47 (Osterbrock & Ferland, 2006). This choice is standard for star-forming galaxies in the literature. Internal reddening was derived from the comparison of the H/H and H/H ratios with their theoretical values. The increase of the Balmer decrement over the intrinsic value due to the extinction can be expressed in terms of the Balmer color excess (Momcheva et al., 2013):
where [(H)/(H)]obs is the observed ratio of the two fluxes. The extinction at the H wavelength is related to the Balmer color excess as following:
where is the assumed extinction curve and the extinction coefficient is defined as . For our targets, we used the Small Magellanic Cloud (SMC) extinction law (Prevot et al., 1984)777This extinction curve is the one that is implemented by default in the tool that we used to determine the ionized gas properties (see Sect. 4.3). Using this curve instead of the more recent extinction measurements by Gordon et al. (2003) does not significantly affect our results. which provides a good representation of a metal-poor interstellar medium.
The observed fluxes of the detected emission lines, ), normalized to (H) and multiplied by 100, the extinction coefficients , and the extinction-corrected intensities normalized to are given in Table 2. The corresponding values for target ID 6, derived from the SDSS spectrum, are provided in the Appendix (Table 5). The uncertainties in the measured fluxes of each line are calculated using the expression from Tresse et al. (1999)
where is the mean standard deviation per pixel of the continuum determined on each side of the emission line, is the B600 spectral dispersion (0.1 nm for 2x binned spectra and 0.05 nm for unbinned spectra), is the number of pixels covered by the line, which we equated to the FWHM of the fitted Gaussian profile, and is the equivalent width of the line.
4.2 BPT diagram
Before deriving the oxygen abundances, we used the Baldwin-Phillips-Terlevich (BPT) diagrams (Baldwin et al., 1981; Veilleux & Osterbrock, 1987) to investigate the source of ionization in these galaxies and determine whether it is driven by star formation or an active galactic nucleus (AGN). AGN in dwarf galaxies are rare to be found, nonetheless an increasing number of studies are discovering intermediate-mass black holes in low-mass systems (Reines et al., 2013; Reines & Volonteri, 2015; Sacchi et al., 2024; Pucha et al., 2024). In Fig. 5 we show the widely used emission-line diagnostic diagram displaying [N ii]/H versus [O iii]/H. The region above the solid line (defined by Kauffmann et al., 2003) indicates that an AGN is the main mechanism of ionization, the region below the dashed line (Kewley et al., 2001) corresponds to a ”pure” star formation ionization mechanism, while the region in between the two curves contain contributions from both AGN and star formation (composite galaxies). The figure shows that the six dwarfs under analysis in this work are typical star-forming galaxies with low [N ii]/H ratios. Similar results are found if we use the [S ii]/H versus [O iii]/H (not shown).
4.3 Oxygen and nitrogen abundances from the direct method
We determined the physical conditions of the star forming regions with the direct method (Pagel et al., 1992; Pilyugin & Thuan, 2005; Osterbrock & Ferland, 2006). Calculations of both the density and temperature of the electron gas were carried out using the Nebular Empirical Analysis Tool (NEAT, Wesson et al., 2012).
1 | 2 | 3 | 4 | 5a | 6 | |
Ne [cm-3] | 10857 | |||||
Te [K] | 18168908 | 166001900 | ||||
O/H | 2.240.02 | 3.29 | ||||
N/H | 0.750.04 | 1.93 | ||||
12+log(O/H) | 7.41 | 7.93 | 7.64 | 7.77 | 7.350.04 | 7.51 |
log(N/O) | -0.93 | -1.39 | -1.25 | -1.25 | -1.470.05 | -1.23 |
a Guseva et al. (2009).
The electron temperature sensitive emission line [Oiii]4363 is detected in all the galaxies (Fig. 4), including the two additional objects with previous spectroscopic observations (ID 5 and 6). Temperatures and densities were derived using an iterative process from the relevant diagnostic lines with NEAT. We used the ratio of the [S ii]67176731 lines to estimate the electron densities, . We found that the ratio is around or above 1.4 for all galaxies, implying that all star-forming regions are in the low-density regime (Osterbrock & Ferland, 2006). Therefore we adopt a constant electron density value of 100 cm-3. The electron temperature ([O iii]) was calculated from the [O iii]4363/(4959 + 5007) emission-line intensity ratio and it was used to derive the abundances of O++ and O+. ranges between 14 000 K and 20000 K. Total oxygen abundances , were obtained from the intensities of the [O ii]73207330 and [O iii] lines. Although not relevant for O/H, additional ionization states that are present in the gas, but not in the optical region of the spectrum, are accounted for with ionization correction factors (ICFs). The ICF used to determine the nitrogen abundance (Kingsburgh & Barlow, 1994) is based on the only bright line detected in the observed wavelength range, [N ii]6584. Uncertainties are propagated from line flux measurements to the derived abundances with a Monte Carlo method implemented by NEAT.
The oxygen abundances for target No. 5 were taken from Guseva et al. (2009), while for galaxy No. 6, we applied the same method as described above, using the [O iii] and [O ii] lines that were observed in the GMOS spectra (see Table 5). We find oxygen abundances varying between 12 + log(O/H) (between 5% and 17% of the solar value), confirming that the galaxies are metal poor. Electron temperatures, electron densities, oxygen and nitrogen abundances are displayed in Table 3.
4.4 The nitrogen-to-oxygen ratio
The nitrogen-to-oxygen ratio, N/O, is considered an indicator of the chemical evolutionary age of a galaxy. While oxygen is mostly formed by primary nucleosynthesis in massive stars ( M⊙) and injected into the ISM by core-collapse supernovae (Nomoto et al., 2013), nitrogen production is more complex and it can have both a primary and secondary origin. Primary nitrogen synthesis can occur in stars with masses around or above 1.3 M⊙ as part of the CNO cycle of oxygen and carbon during hydrogen burning. However stellar evolution models predict that it is mostly produced in intermediate-mass stars (4 - 8 M⊙) and released in the ISM during the asymptotic giant branch phase (AGB, Renzini & Voli, 1981; Woosley & Weaver, 1995; Meynet & Maeder, 2002). At low metallicities, nitrogen is produced from carbon and oxygen created via the helium burning process; in this case, nitrogen behaves like a primary element, its abundance being proportional to that of the other primary heavy elements resulting in a fixed ratio of N/O, independently of the oxygen abundance. Izotov & Thuan (1999) identified that the average value of the N/O ratio of extremely metal-poor galaxies (12 + log(O/H) 7.6) is approximately constant with a very small dispersion (N/O) , the so-called N/O plateau (Fig. 6). This was interpreted as the evidence of nitrogen being produced as a primary element by massive stars at low O/H. If the seed oxygen and carbon of the CNO cycle are produced in previous-generation stars, nitrogen synthesis is called secondary. At high metallicities (12 + log(O/H) ), the ISM metal abundance becomes high enough that the production of secondary nitrogen and N/O increases linearly with oxygen abundance (Tinsley, 1980; Pagel, 1985).
More recent studies found that galaxies with low oxygen abundances show an excess N/O (van Zee & Haynes, 2006; Skillman et al., 2013; Watanabe et al., 2024), indicating that other factors can affect the ratio at low O/H. For example, strong winds carrying nitrogen-enriched gas from post-main sequence hot and massive Wolf-Rayet stars can raise N/O (Walsh & Roy, 1989; Maeder & Meynet, 2001; López-Sánchez et al., 2007; Telles et al., 2014). The increase in the dispersion could also be explained if the galaxy is going through a post-starburst phase: after the production of carbon and primary nitrogen by massive stars during the last major event of star formation, there is a delayed additional contribution of primary nitrogen produced by intermediate-mass stars in the post-starburst phase, unbalanced by the oxygen release of core-collapse supernovae (Garnett, 1990; van Zee & Haynes, 2006; Annibali et al., 2019). Another possibility could be provided by the rapid infall of metal-poor gas from the intergalactic medium that mixes with the ambient metal-rich ISM (Köppen & Hensler, 2005), diluting the oxygen abundance (i.e., it increases H in O/H). As a result, the mixed gas will be characterized by an unusually large N/O ratio for a given O/H (Sánchez Almeida et al., 2016b; Luo et al., 2021). An additional source of dispersion in the N/O versus O/H relation is the fact that oxygen can also be produced – through the third dredge-up in AGB stars – in low-metallicity galaxies, mainly for 12+log(O/H) 7.7, as was proved to happen in the ISM of the irregular dwarf galaxies of the Local Group (Magrini & Gonçalves, 2009).
In Fig. 6, we show the N/O ratios of our targets compared to samples of star-forming galaxies from the literature with different ranges of metallicity (Izotov & Thuan, 1999; Izotov et al., 2006; van Zee & Haynes, 2006; Annibali et al., 2019). This figure shows the N/O plateau identified by Izotov & Thuan (1999) in extremely metal-poor galaxies (orange crosses), although more recent studies, also displayed in the figure, present a larger scatter at low metallicities raising the question about the actual existence of a plateau in this regime (van Zee & Haynes, 2006).
ID | () | () | (SFR)spec | aperture | (SFR)SED |
[erg s-1 cm-2] | [erg s-1 cm-2] | [M⊙ yr-1] | [′′] | [M⊙ yr-1] | |
1 | -13.753 0.002 | -13.66 | -1.04 0.02 | 7.9 | -1.32 0.02 |
2 | -13.728 0.001 | -13.85 | -1.09 0.02 | 7.8 | -1.46 0.02 |
3 | -14.211 0.002 | -14.11 | -2.12 0.03 | 10.5 | -2.29 0.02 |
4 | -13.891 0.001 | -13.53 | -1.99 0.03 | 13.7 | -2.05 0.02 |
5 | -12.936 0.007a | -13.30 | -1.99 0.01 | – | -2.50 0.02 |
6 | -13.814 0.004 | -13.50 | -2.24 0.07 | 3 | -2.38 0.02 |
a Guseva et al. (2009)
We measured the nitrogen abundances and derived the N/O ratios for the four galaxies observed with GMOS and for target N. 6, and we used the values derived in Guseva et al. (2009) for the remaining dwarf (Table 3). We find (N/O) : three out of six dwarfs have typical values given their metallicity range and the scatter observed in other samples, while the remaining three show a higher than expected N/O. Particularly, ID 1 is the galaxy with the largest N/O (log(N/O) = -0.93), which would be more typical of a galaxy with a ten times larger metallicity. We did not detect the He ii 4686 line, which would suggest the presence of Wolf-Rayet stars in any of the dwarfs.
4.5 H line profile fitting and star formation rates
To derive the star formation rates of our targets, we extracted 1D spectra using larger apertures than those employed to estimate the oxygen abundances, in order to include the entire H emission detected in the GMOS spectra (see Fig. 3). The aperture sizes vary between 78 and 137 (Tab. 4). Figure 7 shows the region of the spectra centered around the H line, obtained for these larger apertures. H fluxes, displayed in Table 4, were obtained modeling the line profiles with PySpecKit.
For most of the targets, a satisfactory fit is achieved using two Gaussian components, except for ID 4, where three Gaussians are needed (Fig. 7). As can be seen in the fourth bottom panel of Fig. 3, this galaxy hosts a brighter component with a wider velocity dispersion (knot 3), and two dimmer ones exhibiting narrower (knots 1 and 2). The kinematical centers of the lower luminosity narrow components are offset from those of the main knot, producing the net effect of an overall asymmetry in the global line profiles. Knot 1 is on the blue side of the spectra (approaching gas), while knot 2 is on the red side of the spectra (receding gas). Despite being separated by only approximately 120 pc, knot 1 appears blueshifted at high velocities compared to knot 2. The offset in wavelength between the two star-forming clumps is , corresponding to a velocity difference of km s-1(bottom-right panel in Fig. 7). This might suggest that knot 1 is kinematically distinct from the rest of the galaxy, and further investigation is needed to better address this issue more thoroughly.
We calculated the H luminosity for each galaxy using the H line fluxes listed in Table 4, the extinction coefficients given in Table 2, and the redshifts provided in Table 1. H luminosities are in the range (LH) erg s-1. The H luminosities were converted into SFR using the calibration of Kennicutt & Evans (2012) for a Kroupa (2001) IMF:
where is in units of erg s-1.
Overall, metal-poor galaxies are found to have starburst-level SFRs; however, our sample shows normal star formation activities with SFRs ranging from 0.01 to 0.1 M⊙ yr-1 and specific star formation rates (sSFRs) around 10-10 yr-1. Figure 8 compares sSFRs and stellar masses of our sample with those estimated by H18 and L23. These two samples cover a wider range of these parameters. Particularly, the selection criteria of H23 are such that their sample is more biased towards higher redshift, more active metal-poor galaxies. The solid line shows the star formation sequence defined by Schiminovich et al. (2007) in the SDSS as a reference for the overall trend between the two parameters.
As can be seen in Fig. 8, lower mass galaxies have higher sSFRs, which are consistent with the “downsizing” scenario (Cowie et al., 1996) predicting that lower mass galaxies are more gas-rich and capable of sustaining significant star formation activity at present epoch. The S-PLUS targets have sSFRs that are compatible with their stellar masses, according to the relation of Schiminovich et al. (2007), but overall they appear less active compare to the other samples of metal-poor dwarfs. The SFRs of our targets may be underestimated since the slits cover only a small fraction of the galactic discs. For comparison we derive the H fluxes from the J0660 narrow-band filter following the three-filter method of Vilella-Rojo et al. (2015). The method uses the and bands to trace the linear continuum and the J0660 filter to identify the line emission. It has already been implemented in similar data from the S-PLUS and J-PLUS surveys (Vilella-Rojo et al., 2015; Lima-Dias et al., 2021; Gondhalekar et al., 2024; Smith Castelli et al., 2024). The method provides a H + [N ii] flux estimate, but from the spectra we can see that the contribution of [N ii] is always negligible compared to the the H flux, thus, we do not apply any correction for [N ii] contamination as suggested in Vilella-Rojo et al. (2015). The fluxes obtained in this way are within at most a factor of 2 compared to those obtained from the GMOS spectra, implying that we are not missing a large fraction of the H emission outside the slits. Comparison with the output of the SED fitting procedure discussed in Sect. 2.2 shows similar results. The SFRs derived from the GMOS spectra agree within a factor of 3 with those estimated from SED fitting (Table 4).
4.6 Mass-metallicity relation
The relation between the stellar mass of galaxies and the metallicity of their interstellar medium, known as the mass–metallicity relation (MZR, Lequeux et al., 1979), is the result of the interplay between the processes that drive and regulate the baryon cycle within galaxies: star formation, stellar feedback, and gas accretion (Davé et al., 2011; Lilly et al., 2013). The relation holds for a broad range of stellar masses, from M⊙ to M⊙. It is steeper at low , then it flattens at a characteristic stellar mass reaching an asymptotic saturation metallicity (Tremonti et al., 2004; Curti et al., 2020). The scatter in the MZR has been observed to correlate with different galaxy properties. Ellison et al. (2008a) reported that galaxies with higher sSFRs or larger half-light radii present systematically lower gas-phase metallicities for a given stellar mass. Mannucci et al. (2010) and Lara-López et al. (2010) showed that the MZR has a secondary dependence on the star formation rate, finding an anticorrelation between SFR and metallicity at fixed stellar mass (i.e. galaxies with higher SFRs tend to have lower metallicities). They proposed that local galaxies are distributed in a three-dimensional space defined by mass, metallicity and SFR (the fundamental metallicity relation). Nonetheless, other studies based on integral field unit observations did not find any significant dependence of the observed MZR on the SFR (Sánchez et al., 2017; Barrera-Ballesteros et al., 2017), and no consensus has been reached so far.
The behaviour of the MZR at low mass has been explained in different ways: either low-mass galaxies have been less efficient at producing metals (Brooks et al., 2007), or gas outflows due to supernova feedback are much more efficient in ejecting metals out of low-mass systems, resulting in a lower effective yield with decreasing stellar mass (Dekel & Silk, 1986; De Rossi et al., 2017). Gonçalves et al. (2014) studied the MZR at the low-mass regime using planetary nebulae as tracers of the oxygen abundance. They found that both dwarf irregular and dwarf spheroidal galaxies in the Local Group follow the same MZR, and that their MZR is also consistent with the global MZR of star-forming galaxies (see also Figs. 6 and 7 in Gonçalves, 2019).
Having obtained the oxygen abundances, in Fig. 9 we investigate the MZR for these galaxies and infer whether they follow the general trend observed in galaxies of the Local Volume Legacy (LVL) survey (Berg et al., 2012). Our sample of galaxies (hexagon symbols), as well as other objects from the comparison samples, appear to be offset from the main relation defined by the LVL (black line in Fig. 9). The two most metal-poor galaxies of our sample, ID 1 and 5, are the most extreme outliers. Particularly, the estimated oxygen abundance of ID 1 ( 1/20th the solar value) would be typical of a galaxy with a stellar mass of 1.6 106 M⊙, almost 200 times smaller than the value we measured (Tab. 1, Fig. 9).
The presence of metal-poor outliers at a given stellar mass is in general attributed to external events. This can happen if star formation occurs in pockets of pristine gas accreted from the intergalactic medium (Sánchez Almeida et al., 2016a; De Rossi et al., 2017; Ju et al., 2022; del Valle-Espinosa et al., 2023), or if it is triggered by a merger/interaction of a close galaxy pairs (Ellison et al., 2008b; Lelli et al., 2012; Grossi et al., 2020). In these scenarios, the metallicity of the ISM will be diluted and the galaxy would appear as an outlier in the MZR relation as well as in the luminosity-metallicity relation (LZR). Such an effect would be even stronger in the LZR (Ekta & Chengalur, 2010; McQuinn et al., 2020) due to the high luminosity of young massive stars and the contribution of strong emission lines, making the star-forming dwarfs brighter in the band when compared to galaxies of similar mass with a lower star formation activity.
Outflows due to supernova feedback ejecting enriched gas from the star-forming regions could generate metal losses in low-mass systems (McQuinn et al., 2015, 2019). However, given the low star formation rates observed, it is less likely that strong outflows due to stellar feedback may be causing the observed metal deficiency (Grossi et al., 2020).
To better illustrate the discrepancy in the MZR and the LZR we show in Fig. 10 the luminosity and stellar mass offsets of the three samples from Fig. 9. We define the offsets (, ) as the horizontal distance of each object from the MZR and LZR defined by the relation obtained for the LVL galaxies888We calculate magnitudes using the empirical color transformation of Cook et al. (2014): = + 1.27 + 0.16. () represents the logarithmic difference between the observed stellar mass ( absolute magnitude) of the galaxy and the expected stellar mass ( absolute magnitude) for a given metallicity, if it followed the LVL mass (luminosity)-metallicity relation (see L23 for details). Several galaxies from the L23 and H18 samples exhibit small or negative scatter from the MZR, however, almost half of the galaxies of both samples are located on the upper right side of the plot, with the S-PLUS galaxies being among the most extreme outliers in both relations. DR4_3_SPLUS-s27s07_0011404 (ID 1), the most metal-poor galaxy among the targets observed with GMOS (see Table 3), is the most extreme outlier, lying at the top-right corner of Fig. 10. This galaxy and target No. 5, seem to have a fainter nearby companion at a projected distances of 3 and 2 kpc, respectively (see Fig. 1). However, while in the first case the two systems appear to be interacting, in the second case it is not clear from the Legacy image whether the faint galaxy is actually a companion or just a background object.
A detailed study of the environment of the sample of dwarf galaxies selected in this work is beyond the scope of this paper. Preliminary analysis, using the density catalogue of galaxies in the S-PLUS survey based on the -th nearest neighbor method (Lopes-Silva et al., in prep.), indicate that these systems mostly reside in low-density environments. This will, however, be discussed in more detail in a future work (Gonçalves et al., in prep.).
4.7 Optical morphologies
As a further test, we looked for evidence of perturbations in the optical morphologies that may give hints on galaxy-galaxy interactions or mergers. We analyzed the band images of all 47 galaxies of our sample (see Sect. 2.1) which were taken from the DESI Legacy Imaging Survey. Following Krabbe et al. (2024) we built the diagnostic diagrams involving combinations of non-parametric morphological parameters such as the Gini coefficient (), the second moment of the brightest pixels of a galaxy containing 20 per cent of the total flux (; Lotz et al., 2004), and the concentration–asymmetry–smoothness system (CAS; see Conselice, 2003, for details), which were obtained using the astromorphlib software developed by Hernandez-Jimenez & Krabbe (2022). These diagnostics have been extensively applied to quantify galaxy morphologies and to identify systems with signatures of recent or ongoing mergers. Particularly, the position of a galaxy in the plane allows us to separate major mergers from non-interacting galaxies. Mergers occupy the region defined by the following relation, = 0.14 (Lotz et al., 2008, solid line in the top panel of Fig. 11). Recently, Krabbe et al. (2024) investigated the correlation between CAS morphological parameters to identify galaxies with perturbed morphologies that could be tracing either tidal interactions or ram pressure stripping effects. They identified the diagram as the best diagnostic to separate perturbed from unperturbed galaxies and the separation between these two regions occur at = 9.5 4.85 (solid line in the bottom panel of Fig. 11).
Half of the dwarf galaxies selected with S-PLUS appear to lie in the merger/interacting regions defined by the Gini- and the diagrams, including two targets of this work. ID 2 exhibits an extended, low surface brightness tail, while ID 6 shows a clumpy morphology with star-forming knots distributed throughout its disc, and a high Asymmetry ( = 0.36). Notably, its Legacy image does not reveal features such as tidal tails or stellar streams (Fig. 12). Visual inspection of the Legacy images for other galaxies with high asymmetry and Gini- values reveals elongated and disturbed morphologies, including tails, rings, and possible companions, all indicative of recent or ongoing interactions (Fig. 13).
This analysis may suggest that the low-metallicity of these systems is related to a dwarf-dwarf merger or interaction. However, this result would not rule out that the remaining galaxies might be experiencing minor merger events. Hoyos et al. (2012) report that in late minor mergers in which the less massive galaxies have been almost entirely dissolved, the application of these diagnostics is less effective. Another possibility within the dwarf-dwarf galaxy interaction scenario, is that the stellar discs have not yet been significantly disturbed by the interaction, while the external gas component of the less massive companion has been already stripped and accreted by the most massive galaxy (Werk et al., 2011), as it occurs in the case of IZw18 (Lelli et al., 2012). Lastly, if the star formation process is triggered by accretion from the intergalactic medium, the stellar disc will not be perturbed, but the galaxies will appear as metal-poor outliers in the mass-metallicity relation (Sánchez Almeida et al., 2016a; Zenocratti et al., 2022).
5 Summary and conclusions
We selected a sample of 47 star-forming dwarfs from the S-PLUS survey that are metal-poor galaxy candidates based on their colors, and SED fitting using the whole set of 12 filters provided by S-PLUS. We started a spectroscopic follow-up campaign with the Gemini-S telescope to confirm their low metal abundances and to verify the photometric selection method that we applied. We observed four galaxies and derived O/H using the direct method. Oxygen abundances for two additional systems were derived from previous studies and spectroscopic measurements reported in the literature. We obtained the following results:
1. All the objects with spectroscopic follow-up were found to be metal-poor galaxies, with 12 + log(O/H) , i.e., between 5% and 17% of the solar abundance, confirming the effectiveness of our method based on color cuts and multi-band SED fitting.
2. We estimated nitrogen abundances and nitrogen-to-oxygen ratios. Half of the galaxies have N/O ratios compatible with the observed metallicity, while three dwarfs show excess N/O for a given O/H. Particularly, ID 1, the most metal-deficient target of the sample, exhibits a N/O more typical of a galaxy with ten times higher metallicity.
3. We derived the MZR relation and compared our targets with other samples of dwarf galaxies selected in the SDSS and DES surveys. The S-PLUS galaxies are outliers of the MZR, implying that they are metal-deficient for their stellar masses. For example, the oxygen abundance of ID 1 (approximately 1/20th of the solar value) is typical of a galaxy with a stellar mass nearly 200 times smaller than what we measured.
4. We analyzed the optical morphology of all the 47 metal-poor dwarf galaxy candidates. Half of the sample appears to have perturbed morphologies that might be related to a recent major merger event, according to their location in the Gini- and diagnostic diagrams. Among the six galaxies with a spectroscopic follow-up, only two – ID 2 and 6 – appear to have disturbed optical morphologies according to these diagnostics. Two other systems – ID 1 and 5 – present faint companions in the optical images within a few kpc project distances.
5. The analysis of the H emission line profile of target ID 4 highlights the presence of a star-forming knot with a velocity difference of v 370 km s-1 from the closest H ii region at approximately 120 pc. Such a velocity difference is significantly larger than the typical rotation velocity gradient of a dwarf galaxy. This might imply that the knot is kinematically distinct from the rest of the galaxy; however, further analysis is required to confirm this possibility.
We suggest that an external event may have occurred in these galaxies, triggering star formation and lowering the oxygen abundance. Such an event might be associated, at least in a few cases, with an interaction/merger with a dwarf galaxy companion. Given the low star formation rates observed, it is less likely that strong outflows due to stellar feedback may be causing the observed metal deficiency. Integral field unit observations of the ionized gas, as well as high resolution maps of the atomic gas, will be required to confirm the dwarf-dwarf galaxy interaction scenario.
In future work, we plan to extend the spectroscopic analysis to the other galaxies of the sample in order to assemble a statistically significant set of metal-poor galaxies in the southern hemisphere and to investigate more thoroughly the processes that affect and regulate the metallicity of low-mass systems.
Appendix A Emission line fluxes of DR4_3_SPLUS-n03s22_0021537
4340 H | 48.86 2.57 | 50.45 2.13 |
4363 [O iii] | 6.50 1.47 | 6.71 1.20 |
4861 H | 100.00 2.32 | 100.00 2.70 |
4959 [O iii] | 102.62 3.33 | 102.06 2.74 |
5007 [O iii] | 297.76 7.91 | 295.40 6.48 |
5876 He i | 11.20 1.07 | 10.70 0.88 |
6300 [O i] | 2.78 0.45 | 2.61 0.37 |
6312 [S iii] | 1.85 0.47 | 1.74 0.38 |
6363 [O i] | 1.59 0.41 | 1.50 0.33 |
6563 H | 320.86 7.96 | 300.46 6.19 |
6584 [N ii] | 7.69 0.62 | 7.19 0.50 |
6678 He i | 2.96 0.43 | 2.76 0.35 |
6716 [S ii] | 20.30 0.85 | 18.93 0.68 |
6731 [S ii] | 14.32 0.70 | 13.35 0.57 |
7065 He i | 2.03 0.40 | 1.88 0.32 |
7136 [Ar iii] | 6.58 0.77 | 6.07 0.63 |
7320 [O ii] | 2.96 0.41 | 2.72 0.33 |
7330 [O ii] | 2.28 0.39 | 2.09 0.32 |
4.79 | ||
mean |